3.6 g NaCI is dissolved in 340 grams of water. Calculate the mole fraction of both the solute and solvent?
1
Expert's answer
2020-04-08T13:32:36-0400

Answer:

NNaCl=nNaCl/(nNaCl+nH2O)-mole fraction of NaCl,

NH2O=nH2O/(nNaCl+nH2O)-mole fraction of H2O.

nNaCl=mNaCl/MNaCl=3.6g/(23g*mol-1+35.5g*mol-1)=0.06mol

nH2O=mH2O/MH2O=340g/(2*1g*mol-1+16g*mol-1)=18.89mol

NNaCl=0.06/(18.89+0.06)=0.003 or 0.3%

NH2O=18.89/(18.89+0.06)=0.997 or 99.7%

So, the mole fraction of NaCl is 0.3% and the mole fraction of

H2O is 99.7%.